40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14335     Entry time: Fri Dec 7 17:04:18 2018
Author: gautam 
Type: Update 
Category: IOO 
Subject: IMC transmission 
  • Power just before PSL shutter on PSL table = 97 +/- 1 mW. Systematic error unknown.
  • Power from IFO REFL on AP table = 40 +/- 1 mW. Systematic error unknown.

Both were measured using the FieldMate power meter. I was hesitant to use the Ophir power meter as there is a label on it that warns against exceeding 100 mW. I can't find anything in the elog/wiki about the measured inesrtion loss / isolation of the input faraday, but this seems like a pretty low amount of light to get back from PRM. The IMC visibility using the MC_REFL DC values is ~87%. Assuming perfect transmission of the 87% of the 97mW that's coupled into the IMC, and assuming a further 5% loss between the Faraday rejected port and the AP table, the Faraday insertion loss would be ~30%. Realistically, the IMC transmission is lower. There is also some part of the light picked off for IPPOS. Judging by the shape of the REFL spot on the camera, it doesn't look clipped to me.

Either way, seems like we are only getting ~half of the 1W we send in on the back of PRM. So maybe it's worth it to investigate the situation in the IOO chamber during this vent.


c1pslc1susaux,c1iool0,caux  crates were keyed. Also, the physical shutter on the PSL NPRO, which was closed last Monday for the Sundance crew filming, was opened and the PMC was locked. PMC remains locked, but there is no light going into the IMC.

ELOG V3.1.3-